OpenStudy (anonymous):

solve the equation 6x^3=-1296

OpenStudy (anonymous):

divide both sides by 6, then cube root to get the ans

Directrix (directrix):

6x^3=-1296 x^3 = - 216 x^3 = (-6) ^ 3 x = ? @sonja_lee2

OpenStudy (anonymous):

is it -6?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

That is correct.